Serena is making an experiment. For that, she needs 20 grams of a 52% solution of salt. She has two large bottles of salt water: one with 40% and the other with 70% of salt in them. How much of each must she use to make the solution she needs?
Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

x grams of 40% and 20-x grams of 70%
0.40x+0.70(20-x)=0.52*20
0.40x-0.70x+14=10.40
-0.30x=-3.60
x=12 grams of 40% or 4.8 gms pure and
20-x=8 grams of 70% or 5.6 gms pure.

Answer by ikleyn(52786)

.

May be, it will be interesting to you to know that 52% solution of salt DOES NOT EXIST in the nature.

The highest possible concentration of salt in water is about 28%.

Then the solution becomes SATURATED, and can not dilute more salt.

See this Wikipedia article https://en.wikipedia.org/wiki/Saline_water